OpenStudy (anonymous):

Convert 40/6 to a mixed number in simplest form

OpenStudy (anonymous):

please type how you did the problem please..... *W*

OpenStudy (anonymous):

6 2/3 6*6 = 36 + 4 =40 6 4/6 /2 on each side and get 6 2/3

hero (hero):

Keep subtracting factors of one until you finally get a fraction that is not improper. Then count the number of times you subtracted. The number of times subtracted will be your whole number. Then combine that with the remaining fraction.

OpenStudy (anonymous):

thank u! -_-

OpenStudy (anonymous):

Welcome

hero (hero):

 I subtracted 6 times, and the remaining fraction is 2/3, therefore the mixed fraction is 6 2/3